# OneVL: One-Step Latent Reasoning and Planning with Vision-Language Explanation

### 04 May 2026 v2

@@ 1 Introduction
A growing line of work pursues exactly this direction, replacing explicit reasoning tokens with compact latent representations. COCONUT [40] introduced curriculum learning over latent thought tokens, progressively replacing discrete reasoning steps with continuous vectors. CODI [86] extended this with self-distillation, training a student to mimic a teacher’s CoT behavior in latent space. SIM-CoT [105] attached a separate text-decoding auxiliary decoder to enable direct text supervision during latent training. However, adapting these methods to VLA-based driving reveals critical shortcomings. COCONUT, CODI, and SIM-CoT were designed for language-only reasoning and make no use of the rich visual structure that defines driving scenes. As a result, their purely linguistic latents prove insufficient for the multimodal reasoning demanded by trajectory prediction, and as Figure 1 shows, every existing latent CoT method underperforms explicit CoT across all benchmarks.
@@
Figure 2: Comparison of three CoT paradigms. (a) Explicit CoT: the model generates a full chain of discrete reasoning tokens before the answer. (b) Implicit CoT: reasoning is compressed into a small number of opaque latent vectors $\mathcal{Z}$. (c) OneVL (Ours): two types of latent tokens ($\mathcal{Z}_{v}$, red) and language ($\mathcal{Z}_{l}$, salmon); during training, dual auxiliary decoders decode these into future-frame visual tokens and CoT text, respectively, providing rich text and world model supervision. During inference, the decoders are discarded, and the latent tokens are prefilled into the prompt context, matching the speed of answer-only prediction while keeping the interpretability of (a) in both vision and language.
@@ 2.2 VLM and VLA for Autonomous Driving
Beyond the foundational VLM works and CoT-augmented driving models discussed in Section˜1, a parallel line of research has focused on establishing richer evaluation and supervision signals for language-grounded driving [107, 44]. MapLM [9] introduced a large-scale benchmark specifically targeting map and traffic scene understanding, probing whether VLMs can parse structured road topology from sensor data. Ding et al. [22] augmented VLMs with bird’s-eye-view feature injection, enabling holistic scene understanding that fuses camera and top-down spatial context within a single multimodal model. Complementary efforts have explored corner-case evaluation [12] and risk localization [81] to stress-test VLM reasoning under rare or safety-critical scenarios.
@@ 2.3 World Modeling for Autonomous Driving
The concept of the world model originates from model-based reinforcement learning, where it seeks to emulate human cognitive processes and predict the effects of actions on environmental evolution [39, 38, 54], particularly in 3D and 4D spaces [56, 7, 63, 110, 55, 57]. To further enhance spatial reasoning, several approaches incorporate advanced perception frameworks [100, 116, 117, 108, 109] to achieve a more robust understanding of 3D environments. With advances in video generation and the introduction of the Joint Embedding Predictive Architecture by Assran et al. [2], the scope and applications of world models have broadened considerably [43, JEPA2, 92, 24, 45, 23]. In autonomous driving, world models are typically applied to three ends: data generation, closed-loop evaluation, and representation learning [28, 35, 56, 64, 93].
@@ 4.1 Preliminary: Visual Auxiliary Decoder Self-Supervised Pretraining
Before integrating the visual auxiliary decoder into the full OneVL pipeline, we first pretrain it independently as a future-frame generator. The intuition is straightforward. Asking the decoder to immediately predict future frames conditioned on latent tokens that carry no information yet (early in training) is an ill-posed task that impedes learning. Instead, we first train the decoder with a strong unconditional prior—given the current-frame ViT embeddings, predict what the scene will look like at next two timestamps—before introducing the latent conditioning signal. This preliminary stage is conceptually analogous to self-supervised video prediction where the decoder learns purely from visual observations without any reasoning supervision.
@@ 4.2 Stage 0: Main Model Warmup
The fundamental prerequisite for auxiliary decoders to provide meaningful supervision is that the main model’s latent tokens (i.e., <|latent-vis|> or <|latent|>) carry information that is semantically aligned with reasoning content. Without targeted warmup, these tokens would not produce meaningful hidden states that auxiliary decoders can decode.
@@
  • Predict accurate trajectories: The main CE loss $\mathcal{L}_{\text{CE}}$ over the latent tokens and trajectory answer tokens ensures the model develops strong base prediction capability.
@@ 5.1 Datasets
We evaluate OneVL on four complementary benchmarks: NAVSIM [19], ROADWork [33], Impromptu [17], and APR1. These datasets are chosen because they have been shown to be effective in settings where CoT reasoning is employed, or because they provide sufficient labels to construct explicit reasoning traces [78, 77, 17].
@@
  • APR1 introduces the Chain of Causation (CoC) annotations, featuring decision-grounded reasoning traces aligned with complex driving behaviors to enhance the interpretability and generalization of VLA models.
@@
A key challenge in training OneVL is obtaining high-quality chain-of-thought reasoning annotations paired with each driving scenario. On NAVSIM, we leverage the CoT annotations released by AdaThinkDrive [78], the previous state-of-the-art method. These annotations provide natural-language reasoning traces that cover scene interpretation (such as lane boundaries), critical object analysis (including vehicles and pedestrians), and the final driving intent. They are synthesized by a VLM that converts raw detection labels (e.g., objects and lanes) into reasoning sequences, and serve as the supervision target for the language auxiliary decoder ($\mathcal{L}_{\text{l}}$). On ROADWork, we construct CoT annotations using a similar in-house pipeline. On Impromptu, we build CoT annotations from the original dataset’s Q&A pairs, augmenting the data with explicit decision and root-cause labels for corner-case trajectory prediction. On APR1, we use the released checkpoint to predict the CoC labels for all training examples. These annotations supervise the language auxiliary decoder, enabling the model to learn robust reasoning and decision-making logic in unstructured driving scenarios. Further details are provided in Appendix 8.3.
@@ 5.2 Experimental Setups
On NAVSIM, all methods are evaluated using the Predictive Driver Model (PDM) score, a composite metric that jointly assesses trajectory safety, comfort, and progress. On ROADWork, we report ADE (Average Displacement Error) and FDE (Final Displacement Error) to measure waypoint accuracy. On Impromptu, in addition to ADE and FDE, we also report the trajectory prediction L2 error over the first four seconds, following the protocol of the original paper. On APR1, we report ADE and FDE as well. Across all methods, we additionally report the average inference latency.
@@
On ROADWork, we compare against YNet [33] and on Impromptu, we compare against the Impromptu VLA [17]. For Impromptu VLA, we report the result from our own replication. We also include a result obtained using the provided model checkpoint, which is shown in Table 10. On APR1, we compare against the Cosmos-Reason, a flow-matching based VLA model. AR-based baselines are trained for 2 epochs with a learning rate of $4\times 10^{-5}$ and batch size 64. Latent CoT baselines are trained for 6 epochs with a learning rate of $4\times 10^{-5}$ and batch size 64. The results of previous state-of-the-art methods, where not explicitly stated otherwise, are taken directly from the literature.
@@ 5.3 Main Results
Table 1: Performance comparisons on the NAVSIM benchmark. PDM-score (higher is better) and average inference latency (lower is better) for all methods. indicates the result is derived from the corresponding paper. For OneVL, we only count the parameters of the main VLM, as the auxiliary decoders are discarded during inference. The same applies to all subsequent models.
@@
Table 2: Performance comparisons on the ROADWork benchmark. ADE and FDE (pixels; lower is better), latency (lower is better). indicates the result is derived from the corresponding paper.
@@
$\circ$ Cosmos-Reason 10B 2.86 7.42 Language
@@
Table 5: Performance comparisons on the APR1 benchmark. ADE and FDE (meters; lower is better), latency (lower is better).
Table 5.3 presents the NAVSIM comparison results. Table 5.3 reports ADE/FDE on ROADWork. Table 5.3 and Table 5.3 reports the score on Impromptu. Table 5.3 reports the overall results on APR1. Figure 1 provides a visual overview of all the results. We can make the following observations from the results.
@@
Specifically, OneVL achieves 88.84 PDM-score, surpassing the previous supervised finetuned SOTA AdaThinkDrive (86.20) and LaST-VLA (87.30) by +2.64 and +1.54, respectively. On ROADWork, OneVL also achieves 12.49 ADE and 28.80 FDE, significantly surpassing the previous SOTA YNet (22.68/80.78), respectively. The same pattern is observed on the Impromptu and APR1 datasets. Besides, OneVL outperforms all AR-based, latent CoT, and notably, explicit CoT baselines. This consistent superiority demonstrates the effectiveness of our multimodal auxiliary supervision approach in advancing planning tasks. On the APR1 dataset, OneVL underperforms in FDE with a slight margin (7.53 vs. 7.42). This is expected, as Cosmos-Reason uses RL to further enhance its capabilities.
@@
On NAVSIM, OneVL with prefill inference achieves 4.46 latency, essentially identical to AR answer-only prediction (4.49s). This validates the core efficiency claim of our approach. By prefilling latent tokens into the prompt context (which is processed in a single parallel prefill pass), we incur negligible additional latency compared to sequential autoregressive generation of explicit reasoning tokens. On ROADWork, Impromptu, and APR1, we observe a similar speed-up effect, where OneVL achieves 4.71s, 4.02s, and 3.23s latency respectively, which are comparable to or even lower than those of answer-only prediction. In all cases, OneVL inference is much faster than the AR CoT+Answer baseline.
@@
All three adapted latent CoT methods perform substantially worse than answer-only AR prediction, except for the APR1 dataset. This is a critical finding that purely linguistic latent CoT approaches that were designed for text-only reasoning tasks do not transfer effectively to the multimodal spatial-temporal domain of autonomous driving trajectory prediction. The compact linguistic latent spaces these methods define are insufficient to support the geometric reasoning required for precise waypoint prediction. From the compression view of intelligence, this failure is fundamental rather than incidental. Concretely, these methods compress language descriptions of the scene, but language is already an abstraction of the physical world—it encodes semantic labels and relationships, not the spatial-temporal causal dynamics that determine future outcomes. In other words, it satisfies the efficiency criterion of the compression principle but not the intelligence criterion. A further likely factor is that these baselines are not trained with OneVL’s staged recipe (Section˜4), so their latent streams may remain poorly aligned with trajectory and multimodal supervision when optimization starts fully coupled.
@@
Comparing AR CoT+Answer (88.29) to AR Answer (87.47) on NAVSIM confirms that explicit reasoning supervision provides meaningful trajectory improvements (+0.80 PDM-score). We can observe similar patterns on ROADWork, Impromptu, and APR1. These results confirm the core motivation for incorporating CoT reasoning into autonomous driving systems and establish the quality ceiling that latent CoT methods aim to approach while reducing latency.
@@ 5.4 Explanation Quality
Beyond trajectory prediction accuracy, OneVL is unique in providing human-interpretable explanations in both language and vision. Figures 4, 5, 6 and 7 provide NAVSIM, ROADWork, Impromptu and APR1 qualitative examples, including a side-by-side trajectory prediction comparison (vs. AR Answer baseline), two future frames from the visual auxiliary decoder, and the text CoT from the language auxiliary decoder. We also quantitatively evaluate the quality of text explanations.
@@
Figure 7: Visualizations of prediction on APR1. Each plot overlays ground-truth (green) and predicted (red) trajectories on the front camera view. More examples can be found in Appendix 8.9.
@@ 5.7 In-Depth Analysis: Where Does the Benefit Come From?
The visual auxiliary decoder contributes +0.87 PDM-score compared to +0.31 for the language auxiliary decoder. This asymmetry reflects the world model role the visual decoder plays. Autonomous driving trajectory prediction is fundamentally a spatial prediction task, and visual token reconstruction, i.e., predicting what the scene looks like 0.5–1.0 seconds later, provides a supervision signal that is inherently aligned with the geometric nature of trajectory prediction. Crucially, future-frame prediction is a world model objective: to minimize reconstruction loss on unseen configurations of agents and road geometry, the visual latent tokens must encode the causal dynamics of the scene, not just its current appearance. Language CoT annotation describes the reasoning process in abstract, symbolic terms; it is valuable for semantic grounding but is one step removed from the physical dynamics that drive trajectory outcomes. The world model decoder thus provides a harder, more causally direct compression target that language supervision alone cannot supply, which is precisely why it contributes a larger performance gain.
@@
  • Absence of staged training: Prior methods are typically optimized without OneVL’s three-stage pipeline (Section˜4), leaving language latents, visual latents, and trajectory prediction poorly aligned at the start of optimization. Our ablation isolates this effect. OneVL w/o visual decoder but with staged training still reaches 87.97—well above the no-staged-training collapse (67.13) and above prior latent CoT baselines—demonstrating that the curriculum alone provides substantial gains independent of the visual decoder. Combining both yields the full 88.84, confirming that staged training and world model supervision are complementary. The curriculum creates a stable latent space that the world model decoder can then push toward genuinely causal representations.

### 08 May 2026 v3

# Xiaomi OneVL: One-Step Latent Reasoning and Planning with Vision-Language Explanation

@@ 1 Introduction
A growing line of work pursues exactly this direction, replacing explicit reasoning tokens with compact latent representations [115]. COCONUT [41] introduced curriculum learning over latent thought tokens, progressively replacing discrete reasoning steps with continuous vectors. CODI [87] extended this with self-distillation, training a student to mimic a teacher’s CoT behavior in latent space. SIM-CoT [106] attached a separate text-decoding auxiliary decoder to enable direct text supervision during latent training. However, adapting these methods to VLA-based driving reveals critical shortcomings. COCONUT, CODI, and SIM-CoT were designed for language-only reasoning and make no use of the rich visual structure that defines driving scenes. As a result, their purely linguistic latents prove insufficient for the multimodal reasoning demanded by trajectory prediction, and as Figure 1 shows, every existing latent CoT method underperforms explicit CoT across all benchmarks.
@@
Figure 2: Comparison of CoT paradigms. (a) Explicit CoT: the model generates a full chain of discrete reasoning tokens before the answer. (b) Implicit CoT: reasoning is compressed into a small number of opaque latent vectors $\mathcal{Z}$. (c) OneVL (Ours): two types of latent tokens ($\mathcal{Z}_{v}$, red) and language ($\mathcal{Z}_{l}$, salmon); during training, dual auxiliary decoders decode these into future-frame visual tokens and CoT text, respectively, providing rich text and world model supervision. During inference, the decoders are discarded, and the latent tokens are prefilled into the prompt context, matching the speed of answer-only prediction while keeping the interpretability of (a) in both vision and language.
@@ 2.2 VLM and VLA for Autonomous Driving
Beyond the foundational VLM works and CoT-augmented driving models discussed in Section˜1, a parallel line of research has focused on establishing richer evaluation and supervision signals for language-grounded driving [108, 45, 18]. MapLM [9] introduced a large-scale benchmark specifically targeting map and traffic scene understanding, probing whether VLMs can parse structured road topology from sensor data. Ding et al. [23] augmented VLMs with bird’s-eye-view feature injection, enabling holistic scene understanding that fuses camera and top-down spatial context within a single multimodal model. Complementary efforts have explored corner-case evaluation [12] and risk localization [82] to stress-test VLM reasoning under rare or safety-critical scenarios.
@@ 2.3 World Modeling for Autonomous Driving
The concept of the world model originates from model-based reinforcement learning, where it seeks to emulate human cognitive processes and predict the effects of actions on environmental evolution [40, 39, 55], particularly in 3D and 4D spaces [57, 7, 64, 111, 56, 58]. To further enhance spatial reasoning, several approaches incorporate advanced perception frameworks [101, 118, 119, 109, 110] to achieve a more robust understanding of 3D environments. With advances in video generation and the introduction of the Joint Embedding Predictive Architecture by Assran et al. [2], the scope and applications of world models have broadened considerably [44, 93, 25, 46, 24]. In autonomous driving, world models are typically applied to three ends: data generation, closed-loop evaluation, and representation learning [29, 36, 57, 65, 94].
@@ 4.1 Preliminary: Visual Auxiliary Decoder Self-Supervised Pretraining
Before integrating the visual auxiliary decoder into the full OneVL pipeline, we first pretrain it independently as a future-frame generator. The intuition is straightforward. Asking the decoder to immediately predict future frames conditioned on latent tokens that carry no information yet (early in training) is an ill-posed task that impedes learning. Instead, we first train the decoder with a strong unconditional prior—given the current-frame ViT embeddings, predict what the scene will look like at the next two timestamps—before introducing the latent conditioning signal. This preliminary stage is conceptually analogous to self-supervised video prediction, where the decoder learns purely from visual observations without any reasoning supervision.
@@ 4.2 Stage 0: Main Model Warmup
The fundamental prerequisite for auxiliary decoders to provide meaningful supervision is that the main model’s latent tokens (i.e., <|latent-vis|> or <|latent|>) carry information that is semantically aligned with reasoning content. Without a targeted warmup, these tokens would not produce meaningful hidden states that auxiliary decoders can decode.
@@
  • Predict accurate trajectories: The main CE loss $\mathcal{L}_{\text{CE}}$ over the latent tokens and trajectory answer tokens ensures the model develops a strong base prediction capability.
@@ 5.1 Datasets
We evaluate OneVL on four complementary benchmarks: NAVSIM [20], ROADWork [34], Impromptu [17], and Alpamayo-R1. These datasets are chosen because they have been shown to be effective in settings where CoT reasoning is employed, or because they provide sufficient labels to construct explicit reasoning traces [78, 80, 17].
@@
  • Alpamayo-R1 [105] introduces the Chain of Causation (CoC) annotations, featuring decision-grounded reasoning traces aligned with complex driving behaviors to enhance the interpretability and generalization of VLA models.
@@
A key challenge in training OneVL is obtaining high-quality chain-of-thought reasoning annotations paired with each driving scenario. On NAVSIM, we leverage the CoT annotations released by AdaThinkDrive [78], the previous state-of-the-art method. These annotations provide natural-language reasoning traces that cover scene interpretation (such as lane boundaries), critical object analysis (including vehicles and pedestrians), and the final driving intent. They are synthesized by a VLM that converts raw detection labels (e.g., objects and lanes) into reasoning sequences, and serve as the supervision target for the language auxiliary decoder ($\mathcal{L}_{\text{l}}$). On ROADWork, we construct CoT annotations using a similar in-house pipeline. On Impromptu, we build CoT annotations from the original dataset’s Q&A pairs, augmenting the data with explicit decision and root-cause labels for corner-case trajectory prediction. On Alpamayo-R1, we use the released checkpoint to predict the CoC labels for all training examples. These annotations supervise the language auxiliary decoder, enabling the model to learn robust reasoning and decision-making logic in unstructured driving scenarios. Further details are provided in Appendix 8.3.
@@ 5.2 Experimental Setups
On NAVSIM, all methods are evaluated using the Predictive Driver Model (PDM) score, a composite metric that jointly assesses trajectory safety, comfort, and progress. On ROADWork, we report ADE (Average Displacement Error) and FDE (Final Displacement Error) to measure waypoint accuracy. On Impromptu, in addition to ADE and FDE, we also report the trajectory prediction L2 error over the first four seconds, following the protocol of the original paper. On Alpamayo-R1, we report ADE and FDE as well. Across all methods, we additionally report the average inference latency.
@@
On ROADWork, we compare against YNet [34] and on Impromptu, we compare against the Impromptu VLA [17]. For Impromptu VLA, we report the result from our own replication. We also include a result obtained using the provided model checkpoint, which is shown in Table 10. On Alpamayo-R1, we compare against the Cosmos-Reason, a flow-matching-based VLA model. AR-based baselines are trained for 2 epochs with a learning rate of $4\times 10^{-5}$ and batch size 64. Latent CoT baselines are trained for 6 epochs with a learning rate of $4\times 10^{-5}$ and batch size 64. The results of previous state-of-the-art methods, where not explicitly stated otherwise, are taken directly from the literature.
@@ 5.3 Main Results
Table 1: Performance comparisons on the NAVSIM benchmark. PDM-score (higher is better) and average inference latency (lower is better) for all methods. Symbol indicates the result is derived from the corresponding paper. For OneVL, we only count the parameters of the main VLM, as the auxiliary decoders are discarded during inference. The same applies to all subsequent models.
@@
Table 2: Performance comparisons on the ROADWork benchmark. ADE and FDE (pixels; lower is better), latency (lower is better). Symbol indicates the result is derived from the corresponding paper.
@@
Table 5: Performance comparisons on the Alpamayo-R1 benchmark. ADE and FDE (meters; lower is better), latency (lower is better).
@@
$\circ$ Cosmos-Reason 10B $2.86$ $\mathbf{7.42}$ Language
@@
Table 5.3 presents the NAVSIM comparison results. Table 5.3 reports ADE/FDE on ROADWork. Table 5.3 and Table 5.3 report the score on Impromptu. Table 5.3 reports the overall results on Alpamayo-R1. Figure 1 provides a visual overview of all the results. We can make the following observations from the results.
@@
Specifically, OneVL achieves 88.84 PDM-score, surpassing the previous supervised finetuned SOTA AdaThinkDrive (86.20) and LaST-VLA (87.30) by +2.64 and +1.54, respectively. On ROADWork, OneVL also achieves 12.49 ADE and 28.80 FDE, significantly surpassing the previous SOTA YNet (22.68/80.78), respectively. The same pattern is observed on the Impromptu and Alpamayo-R1 datasets. Besides, OneVL outperforms all AR-based, latent CoT, and notably, explicit CoT baselines. This consistent superiority demonstrates the effectiveness of our multimodal auxiliary supervision approach in advancing planning tasks. On the Alpamayo-R1 dataset, OneVL underperforms in FDE with a slight margin (7.53 vs. 7.42). This is expected, as Cosmos-Reason uses RL to further enhance its capabilities.
@@
On NAVSIM, OneVL with prefill inference achieves 4.46 latency, essentially identical to AR answer-only prediction (4.49s). This validates the core efficiency claim of our approach. By prefilling latent tokens into the prompt context (which is processed in a single parallel prefill pass), we incur negligible additional latency compared to sequential autoregressive generation of explicit reasoning tokens. On ROADWork, Impromptu, and Alpamayo-R1, we observe a similar speed-up effect, where OneVL achieves 4.71s, 4.02s, and 3.23s latency respectively, which are comparable to or even lower than those of answer-only prediction. In all cases, OneVL inference is much faster than the AR CoT+Answer baseline.
@@
All three adapted latent CoT methods perform substantially worse than answer-only AR prediction, except for the Alpamayo-R1 dataset. This is a critical finding that purely linguistic latent CoT approaches that were designed for text-only reasoning tasks do not transfer effectively to the multimodal spatial-temporal domain of autonomous driving trajectory prediction. The compact linguistic latent spaces these methods define are insufficient to support the geometric reasoning required for precise waypoint prediction. From the compression view of intelligence, this failure is fundamental rather than incidental. Concretely, these methods compress language descriptions of the scene, but language is already an abstraction of the physical world—it encodes semantic labels and relationships, not the spatial-temporal causal dynamics that determine future outcomes. In other words, it satisfies the efficiency criterion of the compression principle but not the intelligence criterion. A further likely factor is that these baselines are not trained with OneVL’s staged recipe (Section˜4), so their latent streams may remain poorly aligned with trajectory and multimodal supervision when optimization starts fully coupled.
@@
Comparing AR CoT+Answer (88.29) to AR Answer (87.47) on NAVSIM confirms that explicit reasoning supervision provides meaningful trajectory improvements (+0.80 PDM-score). We can observe similar patterns on ROADWork, Impromptu, and Alpamayo-R1. These results confirm the core motivation for incorporating CoT reasoning into autonomous driving systems and establish the quality ceiling that latent CoT methods aim to approach while reducing latency.
@@ 5.4 Explanation Quality
Beyond trajectory prediction accuracy, OneVL is unique in providing human-interpretable explanations in both language and vision. Figures 4, 5, 6 and 7 provide NAVSIM, ROADWork, Impromptu and Alpamayo-R1 qualitative examples, including a side-by-side trajectory prediction comparison (vs. AR Answer baseline), two future frames from the visual auxiliary decoder, and the text CoT from the language auxiliary decoder. We also quantitatively evaluate the quality of text explanations.
@@
Figure 7: Visualizations of prediction on Alpamayo-R1. Each plot overlays ground-truth (green) and predicted (red) trajectories on the front camera view. More examples can be found in Appendix 8.9.
@@ 5.5 Ablation Study
Table 7: Ablation study on each component in OneVL.
@@ 5.7 In-Depth Analysis: Where Does the Benefit Come From?
The visual auxiliary decoder contributes +0.87 PDM-score compared to +0.31 for the language auxiliary decoder. This asymmetry reflects the world model role the visual decoder plays. Autonomous driving trajectory prediction is fundamentally a spatial prediction task, and visual token reconstruction, i.e., predicting what the scene looks like 0.5–1.0 seconds later, provides a supervision signal that is inherently aligned with the geometric nature of trajectory prediction.
Crucially, future-frame prediction is a world model objective: to minimize reconstruction loss on unseen configurations of agents and road geometry, the visual latent tokens must encode the causal dynamics of the scene, not just its current appearance. Language CoT annotation describes the reasoning process in abstract, symbolic terms; it is valuable for semantic grounding but is one step removed from the physical dynamics that drive trajectory outcomes. The world model decoder thus provides a harder, more causally direct compression target that language supervision alone cannot supply, which is precisely why it contributes a larger performance gain.
@@
  • Absence of staged training: Prior methods are typically optimized without OneVL’s three-stage pipeline (Section˜4), leaving language latents, visual latents, and trajectory prediction poorly aligned at the start of optimization. Our ablation isolates this effect. OneVL w/o visual decoder but with staged training still reaches 87.97, well above the no-staged-training collapse (67.13) and above prior latent CoT baselines, demonstrating that the curriculum alone provides substantial gains independent of the visual decoder. Combining both yields the full 88.84, confirming that staged training and world model supervision are complementary. The curriculum creates a stable latent space that the world model decoder can then push toward genuinely causal representations.